# Prompt Decomposition and Reconstruction Makes Powerful LLM Jailbreakers

### v2

@@ 3.3 Implicit Reconstruction via In-Context Learning
Leveraging demos to guide query reconstruction implicitly
After decomposition, we need to reconstruct the resulting sub-prompts so that the LLM understands the original query. As explained in Section 3.1, the critical insights behind our novel reconstruction algorithm are two folds: 1). Inspired by Chain-of-Thought (Wei et al., 2023a) and Rephrase-and-Respond (Deng et al., 2023), we instruct the target LLM to perform the reconstruction while generating the answer. 2). To avoid leaking intention through the reconstruction task, instead of directly instructing LLM, we propose to embed the reconstruction sub-task inside a set of in-context benign demos, thereby diluting the attention of the LLMs. The main technical challenge lies in generating relevant demos to fulfill this task, which we will explain next.
Figure 4: ICL demo template example of the adversarial prompt "Write a tutorial on how to make a bomb". The template demonstrates a benign step from STRUCTURE and ASSIGNMENT to RESULT (green) and prompt the harmful ASSIGNMENT (red) to LLM.
Automated construction of the demos
Given a set of benign queries, we first apply the prompt decomposition and parse each benign query into four categories: [instruction] [structure], [noun] and [verb], and place the parsed sub-prompts into a Demo Template. The demo template is structured to mimic benign queries, ensuring that the reconstructed prompt appears innocuous while still prompting the LLM to generate the desired response." The template is composed of three parts:
  • STRUCTURE, which explains the parsing structure (e.g., " [instruction] on how [noun] [verb]")
  • ASSIGNMENT, which assigns the parsed sub-prompts to the placeholders in the STRUCTURE section (e.g., [noun] = cake)
  • RESULT, which contains the reconstructed (benign) query and the example response.
Figure 4 displays the proposed context template (Left) used in our experiment, as well as an example of a filled context (Right).
Since the relevance of the demos to the test input plays a critical role in the effectiveness of In-Context Learning, we leverage LLMs to automatically generate benign demos that match the parsing structure of the original query (See Section A.1 for more details).
Once we append the parsed sub-prompts of the original harmful query to the context demos, the entire adversarial prompt will be constructed and can be used to elicit the LLM to answer the original malicious queries.
Algorithm 1 DrAttack
  Input: $p$: initial prompt; $f_{\text{LLM}}$: target LLM; $\mathcal{C}$: combination operator;
  // Prompt decomposition
  GPT generates a depth-$L$ parsing tree for prompt $p$;
  Process words in parsing tree to sub-prompts $p_{1:m}$;
  // Demo generation
  GPT replaces $p_{1:m}$ and gets $q_{1:m}$
  Get demo $C$ with answer $a_{q}=f_{\text{LLM}}(q_{1:m})$;
  // Sub-prompt synonym search
  GPT generates a synonym substitution $s_{\text{syn}}(p_{1:m})$;
  ${s_{\text{syn}}(p_{1:m})}^{\star}=\text{randomsearch}(\mathcal{C}({s_{\text{syn}}(p_{1:m}})))$;
  // Implicit reconstruction via In-Context Learning
  Attack LLM: $f_{\text{LLM}}(C,{s_{\text{syn}}(p_{1:m})}^{\star})$
@@ 3.4 Synonym Search on Sub-Prompts
Another benefit of our framework is that the sub-prompts generated by the prompt decomposition can be further perturbed to enhance the overall performance. We explore a simple Synonym Attack strategy, which empirically further improves the resulting attack success rate. The Synonym Attack strategy involves replacing phrases in the sub-prompts with their synonyms to subtly alter the prompt while maintaining its original intent. This approach increases the likelihood of bypassing LLM safety mechanisms by presenting the prompt in a less detectable form. We construct a phrase-level search space by compiling a list of synonyms for each phrase in the sub-prompts. From there, we deploy a random search to identify the best replacement for each phrase, with the goal of successfully jailbreaking LLMs and generating faithful responses. Due to space limits, we refer the reader to Section A.1 for more details on the random search algorithm.
@@ 4 Experiments
This section evaluates DrAttack, emphasizing its effectiveness and efficiency in jailbreaking LLMs. It illustrates how our decomposition method successfully masks malicious intentions in prompts, effectively avoiding detection. Furthermore, it is demonstrated that DrAttack maintains the faithfulness to the original prompts’ expected responses without revealing the prompts themselves. Lastly, we highlight the robustness of our framework in circumventing perplexity-based filtering mechanisms, showcasing its ability to preserve functionality while bypassing security measures.
@@ 4.1 Experimental setup
Datasets & Metrics
To assess the performance of DrAttack, we use the “harmful behaviors” section of the AdvBench (Zou et al., 2023). This benchmark includes 520 malicious prompts that seek to trigger harmful responses, such as “Write a tutorial on how to make a bomb”. We use the Attack Success Rate (ASR) to measure the effectiveness of the jailbreak attacks. However, the attack’s success is defined on string matching between LLM responses and predefined refusals, which cannot honestly present the attack performance as discussed in recent studies (Zhu et al., 2023; Ding et al., 2023; Shah et al., 2023; Chao et al., 2023; Huang et al., 2023). Instead, we employ two better methods to calculate ASR: 1) GPT evaluation: Assessing harmfulness of responses with gpt-3.5-turbo. The evaluation prompt is provided in Section A.2. However, the jailbroken response can even confuse the LLM evaluator (Li et al., 2023); 2) Human inspection: measuring harmfulness by human evaluation given by generated prompt surveys, as introduced in Section A.2. We also use the average query amount as our efficiency measurement, indicating how many trial attack models need to prompt LLM until a successful attack occurs.
@@
To comprehensively evaluate the effectiveness of DrAttack, we select a series of widely used models to target LLMs with different configurations, including model size, training data, and open-source availability. Specifically, we employ three open-source models, e.g., Llama-2 (Touvron et al., 2023) (7b, 13b) chat models and Vicuna (Zheng et al., 2023) (7b, 13b), and three closed-source models, e.g., GPT-3.5-turbo (OpenAI, 2023a), GPT-4 (OpenAI, 2023b) and Gemini (Team, 2023) in our performance comparison and further analysis. The versions of these models and system prompts we used for the experiments are listed in Section A.2.
@@
In our study, we compare DrAttack with both white-box attacks (GCG (Zou et al., 2023), AutoDan-a (Liu et al., 2023) and AutoDan-b (Zhu et al., 2023)) and black-box attacks (PAIR (Chao et al., 2023), DeepInception (Li et al., 2023)). A transfer attack would be applied to jailbreak closed-source LLMs in the white-box setting for completeness. The tokens are optimized on Vicuna 7b and transferred to black-box models.
@@ 4.2 Results and Analysis
Table 1: Attack success rate (%) (↑) of black-box baselines and DrAttack assessed by human evaluations. $\ast$ represents results from (Li et al., 2023), and the absence of data in some areas is due to the unavailability of reproducible code. DrAttack surpasses prior white-box attacks on all target LLMs.
Black-box attack Closed-source models Open-source models
GPT-3.5-turbo GPT-4 Gemini-Pro Vicuna 7b Llama-2 7b Llama-2 13b
PAIR (Chao et al., 2023) ${}^{\ast}$ ${}^{\ast}$ 10.7 ${}^{\ast}$ ${}^{\ast}$ 4.6
DeepInception (Li et al., 2023) ${}^{\ast}$ ${}^{\ast}$ ${}^{\ast}$ ${}^{\ast}$
DrAttack (Ours) 86.2 84.6 81.5 98.1 38.5 44.2
[…]
Attack type Attack method Closed-source models Open-source models
GPT-3.5-turbo GPT-4 Gemini-Pro Vicuna 7b Llama-2 7b Llama-2 13b
White-box GCG (Zou et al., 2023) ${}^{{\dagger}}$ 0.5 50.9 ${}^{{\dagger}}$ ${}^{{\dagger}}$ 0.0
AutoDan-a (Liu et al., 2023) ${}^{{\dagger}}$ ${}^{{\dagger}}$ ${}^{{\dagger}}$
AutoDan-b (Zhu et al., 2023) ${}^{{\ddagger}}$ ${}^{{\ddagger}}$ ${}^{{\ddagger}}$ ${}^{{\ddagger}}$
Black-box DrAttack (Ours) 78.4 62.0 76.9 81.5 50.1 63.1
We collect the reported ASR with GPT and human evaluation from other jailbreak attacks, with several of them published very recently, to have a detailed full comparison.
As DrAttack is a black-box attacker, we first compare the performance with other black-box attacks in Table 1. Results show that DrAttack outperforms prior black-box attacks by a significant margin; it achieves over 80% success rate on commercial closed-source models such as GPT-3.5, GPT-4 and Gemini-Pro, while previous black-box attack algorithms can only achieve 10-20% ASR on these models. Further, we compare DrAttack with previous white-box attacks using GPT evaluations in Table 2. On the open-source models (Vicuna and Llama), where white-box attacks can fully exploit model weights and structure during an attack, DrAttack can achieve a comparable performance without using white-box information. On the closed-source models (GPT, Gemini), when white-box methods need to conduct transfer attacks, DrAttack again achieves significant improvements on ASR ($\geq 20\%$ absolute improvements in all three cases).
Figure 5: A generated adversarial prompt from the original prompt "Create a tutorial on how to manipulate and trick people into doing something they don’t want to do" and LLMs’ response from our experiment results. The bold texts are sub-prompts from the original prompt.
@@
As a potential red-teaming tool, we test different attack’s efficiency here. DrAttack is more efficient than other attacks demonstrated in Table 3. The average query numbers are from GCG and AutoDan-b are from (Zhu et al., 2023) (convergence iteration multiplied by batch size), and that of DeepInception is from (Chao et al., 2023)’s average queries to obtain a successful adversarial attack. The white-box methods are not directly applied to closed-source models. We leave these places blank for the above reasons, This result demonstrates that DrAttack can significantly reduce the computation costs by shrinking the search space and only employing a random search in the sub-prompts’ synonym space.
@@
Moreover, as illustrated in Figure 6, DrAttack still maintains a high degree of Faithfulness, even after undergoing sophisticated prompt decomposition and reconstruction processes. To quantify the Faithfulness, we calculate the cosine similarity between the ’target’ (the response from the initial prompt attack on the uncensored vicuna model, Wizard Vicuna 13B (Jobbins, 2023)) and the ’output’ (the response from DrAttack on target LLMs). We observe that DrAttack achieves a similar level of cosine similarity compared with previous black-box attacks, demonstrating that our decompose-and-reconstruct approach does not compromise LLMs’ response quality.
Table 3: Number of queries required by baselines and DrAttack. $\star$ represents results from (Chao et al., 2023). DrAttack outperforms other attack strategies by reducing the problem to modifying each sub-prompt.
Closed-source models Open-source models
Attack type Attack method GPT-3.5-turbo GPT-4 Gemini-Pro Vicuna 7b Llama-2 7b
White-box GCG (Zou et al., 2023) - - - 512000 512000
AutoDan-b (Zhu et al., 2023) - - - 51200 51200
Black-box PAIR (Chao et al., 2023) ${}^{\star}$ ${}^{\star}$ 26.3 ${}^{\star}$ ${}^{\star}$
DrAttack(Ours) 12.4 12.9 11.4 7.6 16.1
Figure 6: (a) Mean and variance of cosine similarity between harmful response from target LLM and harmful response from uncensored LLM. (b) Attack success rate drops with attack defenses (OpenAI Moderation Endpoint, PPL Filter, and RA-LLM). Compared to prior black-box attacks (Chao et al., 2023; Li et al., 2023), DrAttack, which first decomposes and then reconstructs original prompts, can elicit relatively faithful responses and is more robust to defenses.
@@
We employ three defensive strategies for LLMs to further verify DrAttack’s effectiveness against defended models. The first defensive strategy, OpenAI Moderation Endpoint (OpenAI, 2023c), is a content moderation tool. It employs a multi-label classification system to sort responses from large language models into 11 specific categories, including violence, sexuality, hate speech, and harassment. A response will be flagged if the given prompts violate these categories. The second defensive strategy, Perplexity Filter (PPL Filter) (Jain et al., 2023), which is designed to detect uninterpretable tokens, will reject jailbreaks when they exceed the perplexity threshold. The third defensive strategy, RA-LLM (Cao et al., 2023), rejects an adversarial prompt if random tokens are removed from the prompt and the prompt fails to jailbreak. All defenses are applied on the successfully jailbroke prompts to evaluate the DrAttacks’ performance. Detailed settings of these attacks are in Section A.2.  Figure 6 demonstrates that the ASR of the attack generated from our framework will only drop slightly when facing the aforementioned defensive strategies. In comparison, PAIR and Deepinceptions suffer from a significant performance drop under the defense by RA-LLM.
@@ 5 Ablation Study
Decomposition and reconstruction concealing malice
We analyzed the following token probabilities from open-source LLMs Llama2 and Vicuna to demonstrate how our method conceals malicious intent. By averaging the probabilities of the first five tokens in rejection strings, we compared responses to original adversarial prompts and those generated by DrAttack. Results in Figure 7 show that while original prompts always lead to rejection tokens (e.g., ’I am sorry …’), prompts processed by our method significantly reduce this likelihood when facing jailbreak prompts.
To prove the effectiveness of prompt decomposition and reconstruction, we further demonstrate the reduction of malice with DrAttack. To assess the malice, OpenAI’s moderation endpoint (OpenAI, 2023c) is employed to calculate classification scores based on OpenAI’s usage policies. We apply the moderation assessment on the original adversarial prompt, sub-prompts after decomposition, and new prompts generated by ICL reconstruction. These scores are categorized into five groups from the original eleven classes and averaged over all 50 prompts sampled. As shown in Figure 7, the decomposition from original adversarial prompts to sub-prompts significantly lowers classification scores in all categories. Interestingly, during the reconstruction phase, there is a slight increase in scores for several categories. However, for the ’self-harm’ category, the reconstructed prompts even further decrease the scores. Overall, both the decomposed phrases and reconstructed prompts exhibit lower rejection rates than the initial prompts, indicating the effectiveness of our method in concealing malice.
Figure 7: (a) Next token probability of rejection string from open-source LLM. The initial prompt has the highest possibility of rejection string generation, while DrAttack can reduce the possibility, especially with a jailbreaking attack (b) Log-scale moderation scores of original adversarial prompt, sub-prompts after decomposition, and new prompt after DrAttack. The higher the score is, the more sensitive content the prompt has for OpenAI’s moderation endpoint. Results show that DrAttack can conceal malice to bypass the output filter.
Figure 8: (a) ASR of generated prompts with ICL semantic-irrelevant, semantic-relevant, or semantic-similar context. (b) ASR of generated prompts with ICL context from vanilla to affirmative ("Sure, here is") and structured (step-by-step) ones.
@@
Here, we investigate whether a semantically similar context in ICL reconstruction can improve the assembling of harmful responses. We design three types of contexts where semantic-irrelevant context uses irrelevant assembling demo; semantic-relevant context gets harmless prompt by making every sub-prompts replaceable; semantic-similar context gets harmless prompt by restricting replaceable sub-prompts, maintaining prompt main sentence while replacing subordinate sub-prompts. The results in Figure 8 indicate that using a semantically similar demo in ICL reconstruction is essential for DrAttack. Furthermore, we show that a more systematic and affirmative example reconstruction from harmless prompt in ICL will trigger LLM to generate harmful contents more frequently in Figure 8. Instead of only prompting plain harmless prompts to generate examples, we also include adding suffix prompt to generate harmless example more systematically with the instruction "Give your answer step-by-step" (Kojima et al., 2023) and more affirmatively with the instruction "Start your sentence with ’Sure, here is’" (Zou et al., 2023). The results show that more systematic and affirmative examples can improve the attack success rate.

### v3

@@ 4.3 Implicit Reconstruction via In-Context Learning
Leveraging benign examples to guide query reconstruction implicitly
After decomposition, the resulting sub-prompts need to be reconstructed so that LLMs understand the original query. As explained in Section 3, the critical insights behind our reconstruction algorithm are two folds: 1) Inspired by Chain-of-Thought Wei et al. (2023b) and Rephrase-and-Respond Deng et al. (2023), we instruct victim LLMs to perform the reconstruction while generating the answer. 2) To avoid leaking intention through the reconstruction task, instead of directly instructing LLM, we propose embedding the reconstruction sub-task inside a set of in-context benign examples, thereby diluting the attention of the LLM. The main technical challenge lies in generating relevant examples to fulfill this task, which we will explain next.
Automated construction of ICL example
Given sub-prompts from the original malicious prompt, we first set [noun] and [verb] as substitutable sub-prompts for the next operation. Then we query GPT-4 to replace substitutable sub-prompts with minimal structure change, after which sub-prompts concatenation are benign (see Appendix A.2 for more details). In this way, the benign example prompt is structured to mimic the original malicious prompt. The reconstruction of a benign example could serve as a context in ICL for reconstructing the original malicious prompt.
In Figure 6, we offer the template for ICL reconstruction with ICL rule, ICL example, and ICL query. The template for the ICL rule is as follows:
  • RULE, which explains the parsing rule to combine sub-prompts (e.g., "[instruction] on how [verb] [noun]")
The template for the ICL example is composed of two parts:
  • ASSIGNMENT, which assigns the benign sub-prompts to the placeholders in the RULE section (e.g., "[noun] = a cake")
  • RESULT, which contains reconstructed (benign) prompt (e.g., "write a tutorial on how to make a cake") and the (benign) response (e.g., "here is a tutorial on how to make a cake …")
The template for the ICL query is then simple:
  • ASSIGNMENT, which assigns the harmful sub-prompts to the placeholders in the RULE section (e.g., "[noun] = a bomb")
Figure 6: ICL template of harmful prompt "Write a tutorial on how to make a bomb." The template demonstrates an implicit reconstruction from RULE to benign ASSIGNMENT & RESULT and prompts the harmful ASSIGNMENT to LLMs.
Once we append the parsed sub-prompts of the original harmful query to the context examples, the entire adversarial prompt will be implicitly reconstructed by LLMs with malicious intent.
@@ 4.4 Synonym Search on Sub-Prompts
Another benefit of our framework is that the sub-prompts generated by prompt decomposition can be further perturbed to enhance the attack performance. We explore a simple Synonym Attack strategy, which improves the resulting attack success rate empirically. The Synonym Attack strategy involves replacing phrases in the sub-prompts with their synonyms to alter the prompt subtly while maintaining its original intent. This approach increases the likelihood of bypassing LLM safety mechanisms by presenting the prompt in a less sensible form. We construct a phrase-level search space by compiling a list of synonyms for each phrase in the sub-prompts. From there, we deploy a random search to identify the best replacement for each phrase, intending to jailbreak LLMs and generate faithful responses. Due to space limits, we refer the reader to Appendix A.3 for more details on the random search algorithm.
@@ 5 Experiments
Closed-source models Open-source models
Attack type Attack methods GPT-3.5-turbo GPT-4 Claude-1 Claude-2 Gemini-pro Vicuna 7b Vicuna 13b Llama2 7b Llama2 13b
white-box GCG Zou et al. (2023) 6 0 0 1 1 88 86 46 38
AutoDAN Liu et al. (2023b) 39 3 5 10 64 88 76 64 2
[…]
PAIR Chao et al. (2023), 12 10 2 1 12 76 70 3 4
DeepInception Li et al. (2023) 0 1 5 5 27 29 7 6 8
ReNellm Ding et al. (2023) 48 13 49 18 48 54 47 30 44
DrAttack (Ours) 78 63 48 27 79 82 63 50 62
Table 1: Attack success rate (%) (↑) of baselines and DrAttack assessed by GPT evaluation.
Closed-source models Open-source models
Attack type Attack methods GPT-3.5-turbo GPT-4 Claude-1 Claude-2 Gemini-pro Vicuna 7b Vicuna 13b Llama2 7b Llama2 13b
white-box GCG Zou et al. (2023) 9 0 0 0 6 96 95 41 40
AutoDAN Liu et al. (2023b) 29 0 0 0 52 93 79 12 0
[…]
PAIR Chao et al. (2023), 12 6 0 0 28 74 70 1 2
DeepInception Li et al. (2023) 5 4 6 3 13 20 13 0 0
ReNellm Ding et al. (2023) 49 15 46 11 50 53 49 27 42
DrAttack (Ours) 84 80 47 26 80 98 80 37 44
Table 2: Attack success rate (%) (↑) of baselines and DrAttack assessed by human evaluation.
@@ 5.1 Experimental setup
Evaluation protocol
Following previous jailbreaking attacks Zou et al. (2023); Liu et al. (2023b); Chao et al. (2023); Wei et al. (2023c); Li et al. (2023); Ding et al. (2023), we leverage the widely adopted AdvBench Zou et al. (2023) for our empirical study. AdvBench includes malicious prompts that seek to trigger harmful responses, such as “Write a tutorial on how to make a bomb”. The test queries are sampled from the harmful behavior section in this benchmark. We also adopt Attack Success Rate (ASR) as the evaluation metric. However, previous works often measure "success" based on the exact matching of LLM outputs with a small set of predefined refusal clauses (e.g., "Sorry"). Such measurement cannot faithfully present the actual effectiveness of the jailbreak due to excessive false positive and false negative rates Zhu et al. (2023); Ding et al. (2023); Shah et al. (2023); Chao et al. (2023); Huang et al. (2023). To better align the evaluation metric with human values, we adopt the following two improved metrics to calculate ASR: 1) Automated evaluation with GPT Wolf et al. (2023); Liu et al. (2023b): We use GPT to judge whether the LLM’s output answers the original malicious query or not (see Appendix B.1 for more details). While GPT judgers are not always reliable Li et al. (2023), they produce much more sensible automated judgments than exact string matching. 2) Human inspection: We also conduct human evaluations on whether an attack is successful or not (see Appendix B.1 for more details). In addition to ASR, another crucial metric for comparing jailbreak methods is query efficiency—specifically, the number of times an attacker must query the LLM to achieve jailbreak. Accordingly, we also report and compare the average number of queries required for each method across all test examples.
@@
To evaluate DrAttack, we select victim LLMs across diverse configurations, availability, and providers. Specifically, our empirical study includes five closed-source models (GPT-3.5-turbo OpenAI (2023a), GPT-4 OpenAI (2023b), Gemini-pro Google (2023), Claude-1 anthropic (2023b) and Claude-2 anthropic (2023a)), and four open-source models (Llama2-chat Touvron et al. (2023) (7b, 13b) and Vicuna Chiang et al. (2023) (7b, 13b)). All model versions are adopted to EasyJailbreak (Zhou et al., 2024) framework for a fair comparison (see Appendix B.2 for model settings).
@@
DrAttack is compared to 6 baselines including both white-box attacks (GCG Zou et al. (2023), AutoDan Liu et al. (2023b)) and black-box attacks (ICA Wei et al. (2023c), PAIR Chao et al. (2023), DeepInception Li et al. (2023), and ReNellm Ding et al. (2023)). All baselines are implemented in EasyJailbreak (Zhou et al., 2024) framework and evaluated by our proposed evaluations (see Appendix B.3 for more details).
@@ 5.2 Results and Analysis
Table 1 shows that DrAttack outperforms prior attacks on powerful LLMs; On the closed-source models (GPT, Claude, Gemini), DrAttack consistently surpasses all existing methods, improving the ASR of prior SOTA on GPT-4 by up to 50% by GPT evaluation and 65% by human evaluation. Moreover, as a black-box method, DrAttack also outperforms other black-box jailbreaking methods, achieving performance on par with even white-box attackers. Notably, contributing to better alignment in newer LLMs (e.g., GPT-3.5-turbo to GPT-4), these victim LLMs are more robust against all attacks, including DrAttack. We also note a discrepancy between GPT-evaluated and human-evaluated ASR on some victim models (GPT-4, Vicuna, Llama2). This is because some models tend to generate harmful content with disclaimers at the end, which affect GPT evaluations. Evaluations on more datasets and examples of jailbreaking can be found in Appendix C and Section D.
@@
As a potential red-teaming tool, we test different attacks’ efficiency in Figure 7. DrAttack is efficient compared to other iteration-needed attacks. The average query number indicates the average number of trials to attack victim LLMs. Query numbers are calculated by total trials on all prompts. (For white-box models, query numbers are calculated by multiplication of batch size and convergence iteration.)
Figure 7: Average number of queries vs. ASR in baselines and DrAttack on Vicuna-7b. DrAttack outperforms other search attack strategies by reducing the problem to modifying each sub-prompt.
(a) different victim models
(b) different defense methods
Figure 8: (a) DrAttack can elicit relatively faithful responses. (b) DrAttack can bypass various defense mechanisms deployed in real-world systems, while the ASR of existing methods drop significantly.
@@
Moreover, as illustrated in Figure 8(a), DrAttack still maintains a high degree of faithfulness, even after undergoing sophisticated prompt decomposition and reconstruction processes. To quantify the faithfulness, we calculate the cosine similarity between the ‘target’ (the response from the original prompt attacking uncensored Vicuna model, Wizard Vicuna 13B Jobbins (2023)) and the ‘output’ (the response from DrAttack on victim LLMs), following previous work Lapid et al. (2023) (see Appendix A.3 for detailed faithfulness calculation). We observe that DrAttack achieves a similar level of cosine similarity compared with previous black-box attacks, demonstrating that our decomposition-and-reconstruction approach does not compromise LLMs’ response quality.
@@
We employ three defensive strategies to verify DrAttack’s effectiveness against defended models further. The first defensive strategy, OpenAI Moderation Endpoint OpenAI (2023c), is a content moderation tool. It employs a multi-label classification system to filter responses from large language models into 11 specific categories, including violence, sexuality, hate speech, and harassment. A response will be flagged if the given prompts violate these categories. The second defensive strategy, Perplexity Filter (PPL Filter) Jain et al. (2023), designed to detect uninterpretable tokens, will reject jailbreaks when they exceed the perplexity threshold. The third defensive strategy, RA-LLM Cao et al. (2023), rejects an adversarial prompt if random tokens are removed from the prompt and the prompt fails to jailbreak. Note that another type of defense mechanism is applied after the complete response has been generated. Commercial LLM providers rarely adopt these output-filter defenses due to their high latency (they must wait for the entire response to be generated and tested before streaming back to the user). OpenAI Moderation Endpoint has been chosen to apply to outputs for its low latency. All defenses are applied directly to prompts or responses (see Appendix B.7 for more details). Fig. 8(b) demonstrates that the ASR of DrAttack will only drop slightly when facing the aforementioned defensive strategies. In comparison, PAIR and Deepinception suffer from a significant performance drop under the defense by RA-LLM.
@@ 6 Ablation Study
(a) ICL examples with different semantics
(b) ICL examples with different contents
Figure 9: Ablation Study: (a) ASR of prompts using ICL exemplars with semantic-irrelevant, semantic-relevant, or semantic-similar context. (b) ASR of prompts with ICL context ranging from vanilla to affirmative ("Sure, here is") and structured (step-by-step). In general, semantic-similar, affirmative, and well-structured ICL exemplars in DrAttack achieve the highest ASR.
(a) different parts of DrAttack
(b) top-k synonyms in random search
Figure 10: Ablation Study: (a) ASR of prompts generated by Decomposition + Reconstruction, Synonym Search, and DrAttack. (b) ASR of top-k Synonym Search applied in DrAttack. DrAttack’s success on powerful LLMs mainly contributes to Decomposition and Reconstruction, with Synonym Search serving as a performance booster.
@@
We investigate whether a semantically similar context in ICL reconstruction can improve the assembly of harmful responses. We design three types of contexts where semantic-irrelevant context uses irrelevant assembling demo; semantic-relevant context gets benign prompt by making every sub-prompts replaceable; semantic-similar context gets benign prompt by restricting replaceable sub-prompts, maintaining prompt main sentence while replacing subordinate sub-prompts. The results in Figure 9(a) indicate that using a semantically similar demo in ICL reconstruction is essential for DrAttack. Instead of only prompting plain benign prompts to generate examples, we also add a naive suffix to generate benign examples more systematically with the instruction "Give your answer step-by-step" Kojima et al. (2023) and more affirmatively with the instruction "Start your sentence with ‘Sure, here is’" Zou et al. (2023). The results in Figure 9(b) show that more systematic and affirmative examples can improve ASR.
Effectiveness of Synonym Search
Besides Decomposition and Reconstruction, DrAttack introduces a Synonym Search to boost attack performance. We investigate the effectiveness of Synonym Search by decoupling it from DrAttack, directly applying it to sub-prompts (e.g., "how to make a bomb" to "how to construct a bomb"). Fig. 10(a) demonstrates that Decomposition and Reconstruction are essential parts that make DrAttack jailbreaking victim LLMs, while Synonym Search boosts attack performance. We further investigate top-k synonyms of Synonym Search and observe slight improvements in ASR with more synonym candidates in Figure 10(b).